Bob invested $ 30,00 part at 10% and part at 1%. If the total interest at the end of the year is $1,560 how much did he invest at 10%
Answer by checkley75(3666)

I'LL ASSUME THE AMOUNT TO BE INVESTED IS $30,000 THEN:
.10X+.01(30,000-X)=1,560
.10X+300-.01X=1,560
.09X=1,560-300
.09X=1,260
X=1,260/.09
X=14,000 INVESTED @ 10%.
30,000-14,000=16,000 INVESTED @ 1%
PROOF
.10*14,000+.01*1,600=1,560
1,400+160=1,560
1,560=1,560